OpenStudy (anonymous):

Find the value or r so the line that passes through the pair of points below has the given slope. (5 , r), (4 , 9), m = 1 r =

OpenStudy (anonymous):

use the formula for the slope m=\[\frac{ y _{2}-y_1 }{x_2-x_1 ? }\] where y2, y1 and x2,x1 are x and y coordinates from the points given

OpenStudy (anonymous):

insert values in the formula 1=(r-9)/(5-4) solve for r and you'll get r=1